Use the Laplace transform to solve the given initial value problem.

y′′−4y′−96y=0; y(0)=14, y′(0)= 28


Enclose arguments of functions in parentheses. For example, sin(2x).

y =___________
1
Expert's answer
2020-08-03T18:21:52-0400

Solution:

"y(x)\leftrightarrow Y(p)"

"y'(x)\leftrightarrow pY(p)-y(0)=pY(p)-14"

"y''(x)\leftrightarrow p^2Y(p)-py(0)-y'(0)=p^2Y(p)-14p-28"

We compose an operator equation that corresponds to the given differential equation:

"p^2Y(p)-14p-28-4(pY(p)-14)-96Y(p)=0"

"(p^2-4p-96)Y(p)=14p-28"

"Y(p)=(14p-28)\/(p^2-4p-96)"

"Y(p)=(14p-28)\/((p+8)(p-12))"


"y(x)=Res_{p=-8}(Y(p)e^{px})+Res_{p=12}(Y(p)e^{px})="


"=\lim\limits_{p\rarr-8}\frac{14p-28}{p-12}e^{px}+\lim\limits_{p\rarr12}\frac{14p-28}{p+8}e^{px}="


"=\frac{-140}{-20}e^{-8x}+\frac{140}{20}e^{12x}=7e^{-8x}+7e^{12x}"


Answer: "y(x)=7e^{(-8x)}+7e^{(12x)}"